# Wider or Deeper? Scaling LLM Inference-Time Compute with Adaptive Branching Tree Search

### v3

@@ 4.2 Results
ARC-AGI. Figure 4 (right) shows the performance on ARC-AGI, a particularly challenging benchmark. Following ARC-AGI’s official evaluation protocol, we report Pass@2. Consistent with previous work [13], repeated sampling proves to be a strong baseline in our setup, indicating the importance of broad exploration for this task. While standard MCTS yields only marginal improvements with larger budgets, our AB-MCTS framework achieves performance comparable to repeated sampling. This suggests AB-MCTS’s capability to effectively explore potentially by dynamically widening its search when beneficial. Similar results were observed with DeepSeek-V3, as detailed in Appendix Figure 8.
@@ 4.3 Analysis
Analysis of Search Behavior: Width vs. Depth. To quantitatively analyze how AB-MCTS balances exploration and exploitation, we examined the average depth and the average width at each depth of the generated search trees. Figure 5 shows that AB-MCTS methods tend to generate wider trees compared to standard MCTS. This occurs because AB-MCTS can adaptively decide to explore wider (select the GEN node) from any existing node, unlike standard MCTS. This mechanism allows for more flexible exploration across various tree depths. In addition to this flexibility in exploring wider, as seen in Table 2, AB-MCTS also achieves strong performance on benchmarks where sequential refinement excels, suggesting that AB-MCTS effectively identifies and exploits promising branches by selecting existing child nodes for refinement. This adaptive nature allows it to combine the strengths of exploration and exploitation, resulting in robust performance across diverse benchmarks.
@@
Scaling with Increased Budget. Highly complex problems often require a substantial generation budget to find a correct solution. ARC-AGI is a prime example, where extensive exploration via repeated sampling is known to improve performance even at large budgets as reported in [13]. To investigate the scaling properties of our approach, we extended the experiments on ARC-AGI using DeepSeek-V3 with a larger generation budget up to $5122^{9}=$ 512. As shown in Figure 7, the performance of AB-MCTS continues to improve substantially as the budget increases from $200$ to $500$, while the improvement rate of repeated sampling begins to plateau. Standard MCTS also continues to improve with a larger budget, yet shows a significantly lower success rate compared to the AB-MCTS methods. This performance gap highlights that AB-MCTS is more effective at directing its search towards promising branches within the search tree at large computational scales.
Qualitative Analysis of the Search Trees. Figure 6 and Appendix Figure 11 present example search trees generated by AB-MCTS-M and standard MCTS. These visualizations illustrate more adaptive branching by AB-MCTS-M compared to standard MCTS. This adaptive nature reveals that AB-MCTS-M flexibly balances exploration and exploitation throughout the search process, dynamically allocating budget to explore diverse new candidates ("going wider") and refine promising ones ("going deeper"). Further discussion can be found in Appendix C.3.
@@ A.1.2 Existing Methods for Inference-Time Answer Search
A.2 AB-MCTS-M Details
A.2.1 Background on Mixed Models
@@
A.2.2 Detailed Mixed Model Formulation and Example Code
@@
# Child indices use 0-based indexing; note the difference from index j in Equations (8-9)
@@
A.3 AB-MCTS-A Details: Parameter Update Rules
A.3.1 AB-MCTS-A (Gaussian) Parameter Update Rules
@@
A.3.2 AB-MCTS-A (Beta) Parameter Update Rules
@@
where $B(\cdot\mid\alpha,\beta)$ denotes the Beta distribution. We note that, usually, this update rule is used in conjunction with the Bernoulli trial, but here we directly use Beta distribution to model the score distribution. In practice, this parameter update rule worked well according to our experimental results.
@@ B.1 Tasks and Datasets
MLE-Bench comprises practical machine learning tasks derived from Kaggle competitions. In order to enable fair comparisons [16], we adopt three low-complexity challenges (Nomad2018 Predicting Transparent Conductors, Spooky Author Identification, and Random Acts of Pizza). Each competition’s training data is randomly split into $80$% for training and $20$% for validation. The validation set is used to obtain the scores for each node. We select a node with the highest validation score at a given inference budget and then evaluate it on the hidden test set to get the final result. Following previous research [16], we use the AIDE scaffold for our experiments. Evaluating the generated machine learning models within these competitions is notably resource-intensive, requiring substantial GPU power even to process a single solution candidate. In our experiments, each solution candidate was executed on a single H100 GPU with a time limit of one hour. This computational demand is still considerably higher than for other benchmarks discussed in this work. Consequently, due to these significant costs, comprehensive experimentation across all methods and models on MLE-Bench was prohibitive. We therefore focused our evaluations using GPT-4o for these tasks specifically on AB-MCTS-M.
@@ B.2 AB-MCTS Parameters
  • The average value of score is $0.5$, reflecting a neutral initial assumption regarding answer quality (where 00 indicates the worst and $1$ the best)
@@
We assign the following priors for AB-MCTS-M in Equations 8 and 9:
@@
where $\mathcal{N}_{\text{half}}$ is the half-normal distribution (please refer to Section 3.3 and Appendix A.2). For AB-MCTS-A (Gaussian), we set $\breve{m}=0$, $\breve{\kappa}=1$, $\breve{\nu}=1$, and $\breve{\tau}^{2}=0.1$ in Equations 11 - 16, and for AB-MCTS-A (Beta), we set $\breve{\alpha}=0.5$ and $\breve{\beta}=0.5$ in Equations 17 - 19. As we noted earlier, we chose these parameters in a way that imposes as few assumptions as possible to minimize bias.
@@ C.1 Results with DeepSeek-V3 on Competitive Programming and ARC-AGI
Figure 10: Performance comparison on three MLE-Bench tasks using GPT-4o. Each plot shows performance versus the total generation budget. For Nomad2018 Predicting Transparent Conductors and Spooky Author Identification, lower scores are better (RMSLE and Log Loss, respectively); for Random Acts of Pizza, higher is better (ROC AUC). At each budget, we choose the single solution based on validation-set performance and report its test-set score.

### v5

@@ 4.2 Results
ARC-AGI. Figure 5 (right) shows the performance on ARC-AGI, a particularly challenging benchmark. Following ARC-AGI’s official evaluation protocol, we report Pass@2 (Pass@1 is also reported in Appendix C.6). Consistent with previous work [13], repeated sampling proves to be a strong baseline in our setup, indicating the importance of broad exploration for this task. While standard MCTS yields only marginal improvements with larger budgets, our AB-MCTS framework achieves performance comparable to repeated sampling. This suggests AB-MCTS’s capability to effectively explore potentially by dynamically widening its search when beneficial. Similar results were observed with DeepSeek-V3, as detailed in Appendix Figure 10.
@@ 4.3 Analysis
Analysis of Search Behavior: Width vs. Depth. To quantitatively analyze how AB-MCTS balances exploration and exploitation, we examined the average depth and the average width at each depth of the generated search trees. Figure 5 shows that AB-MCTS methods tend to generate wider trees compared to standard MCTS. This occurs because AB-MCTS can adaptively decide to explore wider (select the GEN node) from any existing node, unlike standard MCTS. This mechanism allows for more flexible exploration across various tree depths (See also Appendix C.7). In addition to this flexibility in exploring wider, as seen in Table 2, AB-MCTS also achieves strong performance on benchmarks where sequential refinement excels, suggesting that AB-MCTS effectively identifies and exploits promising branches by selecting existing child nodes for refinement. This adaptive nature allows it to combine the strengths of exploration and exploitation, resulting in robust performance across diverse benchmarks.
@@
Scaling with Increased Budget. Highly complex problems often require a substantial generation budget to find a correct solution. ARC-AGI is a prime example, where extensive exploration via repeated sampling is known to improve performance even at large budgets as reported in [13]. To investigate the scaling properties of our approach, we extended the experiments on ARC-AGI using DeepSeek-V3 with a larger generation budget up to $2^{9}=512$. As shown in Figure 6, the performance of AB-MCTS continues to improve substantially as the budget increases from $200$ to $500$, while the improvement rate of repeated sampling begins to plateau. Standard MCTS also continues to improve with a larger budget, yet shows a significantly lower success rate compared to the AB-MCTS methods. This performance gap highlights that AB-MCTS is more effective at directing its search towards promising branches within the search tree at large computational scales.
Qualitative Analysis of the Search Trees. Figure 7 and Appendix Figure 11 present example search trees generated by AB-MCTS-M and standard MCTS. These visualizations illustrate more adaptive branching by AB-MCTS-M compared to standard MCTS. This adaptive nature reveals that AB-MCTS-M flexibly balances exploration and exploitation throughout the search process, dynamically allocating budget to explore diverse new candidates (“going wider”) and refine promising ones (“going deeper”). Further discussion can be found in Appendix C.3.
@@ A.1.2 Existing Methods for Inference-Time Answer Search
A.2 Comparison of AB-MCTS and Standard MCTS with Progressive Widening
The progressive widening has parameters $(k,\alpha)$ which bounds the number of branching factors by $kn^{\alpha}$ with node visit count $n$. With these parameters, the rule for whether to branch is pre-determined as a function of the node’s visit count. Crucially, this decision does not use important information gathered during the search, namely, the observed rewards of the expanded nodes. The UCT score is only used to select which child to descend to after the decision has been made not to branch. Furthermore, defining the branching rule as a function of visit count means that the scaling behavior of the tree’s shape and node degrees is pre-determined by the choice of hyperparameters.
In contrast, our approach does not restrict the branching rule based solely on visit counts and hyperparameters. Instead, the branching factor adapts dynamically based on the observed rewards. This is an important requirement for LLM test-time inference scaling, where a tree search that purely goes wide is known to be a strong baseline. To demonstrate the robustness of AB-MCTS, we conducted an experiment to compare AB-MCTS and progressive widening in Appendix C.4.
A.3 AB-MCTS-M Details
A.3.1 Background on Mixed Models
@@
A.3.2 Detailed Mixed Model Formulation and Example Code
@@
# Child indices use 0-based indexing; note the difference from index j in Equations (9-10)
@@
A.4 AB-MCTS-A Details: Parameter Update Rules
A.4.1 AB-MCTS-A (Gaussian) Parameter Update Rules
@@
A.4.2 AB-MCTS-A (Beta) Parameter Update Rules
@@
where $B(\cdot\mid\alpha,\beta)$ denotes the Beta distribution. We note that, usually this update rule is used in conjunction with the Bernoulli trial, but here we directly use Beta distribution to model the score distribution. In practice, this parameter update rule worked well according to our experimental results.
A.5 Walk‑through Examples
We walk through an iteration of AB-MCTS-M and AB-MCTS-A on the example trees in Figures 2 and 3. The process is stochastic due to Thompson sampling; for clarity, we assume specific sampled outcomes.
A.5.1 AB‑MCTS‑M
AB-MCTS-M incrementally builds the search tree by adding one node at a time. In this section, we detail a single iteration of the AB-MCTS-M algorithm, clearly illustrating the sequence of selecting a node to expand, performing the expansion, and backing up the resulting score. For simplicity and concreteness, we assume the current search tree structure is as depicted in Figure 2, and we describe one complete iteration, including selection, expansion, and score backup.
  1. 1.
    ($N\to N_{1}$) At $N$, we compute posterior distributions for its four children, GEN, $N_{1}$, $N_{2}$, and $N_{3}$ under the mixed model, and draw one score from each. If GEN receives the highest sample it is expanded and its score is backed up (Algorithm 1, lines 11–13). Here, we assume that $N_{1}$ attains the highest sampled score, reflecting the exploitation of a child whose posterior peak is comparatively large.
  2. 2.
    ($N_{1}\to N_{1}^{\prime}$) $N_{1}$ has two direct children, $N_{1}^{\prime}\;(r=0.8)$ and $N_{2}^{\prime}\;(r=1.0)$. We compute posteriors for GEN, $N_{1}^{\prime}$, and $N_{2}^{\prime}$, then sample again. Although the subtree $T(N_{2}^{\prime})$ currently contains nodes with higher score as we can see from Figure 2, the finite variance of its posterior ensures that $N_{2}^{\prime}$ is not always selected, and encourages more exploration for under-explored tree regions. Suppose $N_{1}^{\prime}$ is chosen on this iteration.
  3. 3.
    (Expanding $N_{1}^{\prime}$) Because $N_{1}^{\prime}$ is a leaf, we expand it (Algorithm 1, line 10). Assume the newly generated node receives the score $r=0.5$. Since all the leaf nodes have a GEN child, a GEN node is appended to this expanded node as well.
  4. 4.
    (Score backup) The score is propagated upwards from the expanded node toward the root, as in standard MCTS. In the current example, the score is backed up through the node generated at step 3, then through nodes $N_{1}^{\prime}$, $N_{1}$, and $N$. Unlike standard MCTS, AB-MCTS-M maintains individual scores rather than averages. Specifically, the backed-up scores for nodes $N_{1},N_{2},N_{3}$ form distinct observation lists corresponding to each group in the mixed model. Although GEN nodes have no direct observations due to this score backup rule, their posterior distributions share statistical strength with these groups, allowing indirect information sharing and improved estimation accuracy for a score obtained by node expansion. At the next SelectExpansionTarget call, the four posteriors at $N$ have different shapes; specifically, the peak of $N_{1}$’s posterior shifts left due to the lowered expected value. Other posterior distributions are affected as well, e.g., the right‑hand tail of the GEN posterior contracts. Please note that, in AB-MCTS-M, the change of posterior distribution shape cannot be analytically written down, and is calculated by MCMC. Thus, unlike standard MCTS, the score backup step involves appending the new score to a list of scores.
A.5.2 AB‑MCTS‑A
AB-MCTS-A works in a similar manner to AB-MCTS-M, except for the introduction of CONT node and how we perform score backup. To clearly illustrate the algorithm and highlight the difference from AB-MCTS-M, we describe a complete iteration of selection, expansion, and score backup cycle for AB-MCTS-A. Since the only difference between Beta and Gaussian variants is how the score update is reflected in the posterior distribution, here we focus on the qualitative aspect of posterior update and focus on the details for an example tree, Figure 3.
  1. 1.
    ($N\to\text{CONT}$) At the root, we sample from the posteriors of the GEN and CONT children. As detailed in Section 3.4, the GEN posterior is informed by the CONT children nodes $N_{1}\;(0.8)$, $N_{2}\;(0.0)$, $N_{3}\;(0.2)$, whereas the CONT posterior uses the CONT node’s descendant scores excluding $N_{1}$, $N_{2}$ and $N_{3}$, i.e., 0.8, 1.0, 0.3. We assume CONT is selected here.
  2. 2.
    ($\text{CONT}\to N_{1}$) Next we compute posteriors for CONT’s children $N_{1}$, $N_{2}$, and $N_{3}$. Due to the score-backup rule, the posterior distributions are computed from previously expanded nodes; Concretely, the following scores are used for posterior distribution calculation: $N_{1}$: (0.8, 0.8, 1.0), $N_{2}$: (0.0), and $N_{3}$: (0.2, 0.3). Here, we assume $N_{1}$ obtains the highest sample via Thompson sampling.
  3. 3.
    (Expanding $N_{1}$) Again, we perform Thompson sampling between the GEN and CONT children of $N_{1}$. The GEN posterior uses (0.8, 1.0); the CONT posterior falls back to the prior because no generated descendants exist. Here we assume GEN is selected, leading to the expansion of a new node under $N_{1}$’s CONT child. We assume the score $r=0.5$.
  4. 4.
    (Score backup) We backup the score $r=0.5$ as prescribed in Section 3.4. First, the score is backed up to the GEN node which generates the node. Second, it propagates to: (i) GEN node’s ancestor $N_{1}$, and (ii) the CONT node that is $N_{1}$’s parent. Because 0.5 is lower than the existing scores 0.8, 1.0, the posterior peak of $N_{1}$ shifts left, reducing the probability that $N_{1}$ will be chosen again from CONT. Similarly, adding 0.5 to the existing scores (0.8, 1.0, 0.3) lowers the peak of the CONT posterior at $N$, thus decreasing the probability that CONT will be selected at $N$ in later iterations.
A.6 Hyperparameter Sensitivity of AB-MCTS
This section presents a sensitivity analysis of the hyperparameters used in AB-MCTS. As discussed in Appendix B.2, the prior parameters were designed to be non-informative to minimize bias. Since the posterior distributions become increasingly data-driven as the search progresses, we hypothesized that the influence of the initial priors would be limited. Here, we empirically verify this hypothesis through an extensive analysis. We evaluated the sensitivity of AB-MCTS variants to their prior hyperparameters on LiveCodeBench, using GPT-4o with a generation budget of $2^{4}$. Each configuration was run five times ($n=5$) to compute mean and standard deviation of Pass@1. Tables 3 summarize the results for AB-MCTS-M, AB-MCTS-A (Gaussian), and AB-MCTS-A (Beta) under various prior settings. Across all tested ranges, the performance remains stable, indicating low sensitivity to the initial hyperparameter values. This confirms that AB-MCTS is robust to the initialization of prior hyperparameters, and its performance is primarily governed by the data-driven posterior updates during search.
Table 3: Hyperparameter Sensitivity of AB-MCTS. Pass@1 results for each prior setting. All values are averaged over five runs.
AB-MCTS-M
$\breve{m}$ 0.0 0.4 0.5 0.6 1.0
Pass@1 38.4 ± 1.6 37.3 ± 0.4 36.8 ± 1.5 37.5 ± 1.5 37.7 ± 1.3
$\breve{\alpha}$ 0.01 0.1 0.2 0.3 1.0
Pass@1 38.4 ± 1.3 37.7 ± 0.9 36.8 ± 1.5 38.2 ± 2.3 38.6 ± 1.8
$\breve{\tau}$ 0.01 0.1 0.2 0.3 1.0
Pass@1 37.3 ± 2.0 38.2 ± 1.1 37.1 ± 1.2 36.8 ± 1.5 39.5 ± 1.3
AB-MCTS-A (Gaussian)
$\breve{m}$ 0.0 0.1 0.5 1.0
Pass@1 38.0 ± 1.6 37.0 ± 1.4 37.3 ± 1.5 37.7 ± 0.7
$\breve{\kappa}$ 0.001 0.5 1.0 10.0
Pass@1 38.4 ± 1.3 37.3 ± 1.5 38.0 ± 1.6 37.7 ± 0.7
$\breve{\nu}$ 0.001 0.5 1.0 10.0
Pass@1 38.2 ± 1.3 37.7 ± 1.3 38.0 ± 1.6 38.9 ± 1.0
$\breve{\tau}^{2}$ 0.05 0.1 0.2 0.5 1.0
Pass@1 37.5 ± 0.9 38.0 ± 1.6 37.7 ± 1.3 37.9 ± 0.8 37.7 ± 0.7
AB-MCTS-A (Beta)
$\breve{\alpha}$ 0.1 0.4 0.5 0.6 1.0
Pass@1 37.3 ± 1.5 37.0 ± 1.5 37.5 ± 1.3 37.5 ± 1.3 37.7 ± 1.2
$\breve{\beta}$ 0.1 0.4 0.5 0.6 1.0
Pass@1 38.4 ± 1.8 38.4 ± 1.1 37.5 ± 1.3 37.9 ± 0.5 37.9 ± 1.4
@@ B.1 Tasks and Datasets
MLE-Bench comprises practical machine learning tasks derived from Kaggle competitions. In order to enable fair comparisons [16], we adopt three low-complexity challenges (Nomad2018 Predicting Transparent Conductors, Spooky Author Identification, and Random Acts of Pizza). Each competition’s training data is randomly split into 8080% for training and 2020% for validation. The validation set is used to obtain the scores for each node. We select a node with the highest validation score at a given inference budget and then evaluate it on the hidden test set to get the final result. Following previous research [16], we use the AIDE scaffold for our experiments. Evaluating the generated machine learning models within these competitions is notably resource-intensive, requiring substantial GPU power even to process a single solution candidate. In our experiments, each solution candidate was executed on a single H100 GPU with a time limit of one hour. This computational demand is still considerably higher than for other benchmarks discussed in this work. Consequently, due to these significant costs, comprehensive experimentation across all methods and models on MLE-Bench was prohibitive. We therefore focused our evaluations using GPT-4o for these tasks specifically on AB-MCTS-M.
@@ B.2 AB-MCTS Parameters
  • The average value of score is $0.5$, reflecting a neutral initial assumption regarding answer quality (where 0 indicates the worst and $1$ the best)
@@
We assign the following priors for AB-MCTS-M in Equations 9 and 10:
@@
where $\mathcal{N}_{\text{half}}$ is the half-normal distribution (please refer to Section 3.3 and Appendix A.3). For AB-MCTS-A (Gaussian), we set $\breve{m}=0$, $\breve{\kappa}=1$, $\breve{\nu}=1$, and $\breve{\tau}^{2}=0.1$ in Equations 12 - 17, and for AB-MCTS-A (Beta), we set $\breve{\alpha}=0.5$ and $\breve{\beta}=0.5$ in Equations 18 - 20. As we noted earlier, we chose these parameters in a way that imposes as few assumptions as possible to minimize bias.
@@
Figure 10: Performance comparison on three MLE-Bench tasks using GPT-4o. Each plot shows performance versus the total generation budget. For Nomad2018 Predicting Transparent Conductors and Spooky Author Identification, lower scores are better (RMSLE and Log Loss, respectively); for Random Acts of Pizza, higher is better (ROC AUC). At each budget, we choose the single solution based on validation‐set performance and report its test-set score.
Appendix C Additional Experiments and Analysis
@@ C.3 Example search trees generated by each methods on MLE-Bench
Table 4: Comparison between Progressive Widening and AB-MCTS on LiveCodeBench.
($k$, $\alpha$) = (1, 0.45) ($k$, $\alpha$) = (5, 0.5) ($k$, $\alpha$) = (10, 0.55) AB-MCTS-A (Gaussian) AB-MCTS-A (Beta) AB-MCTS-M
Pass@1 48.7 $\pm$0.8 50.7 $\pm$1.3 50.5 $\pm$3.1 48.9 $\pm$2.0 51.8 $\pm$1.4 49.6 $\pm$1.6
C.4 Comparison to Progressive Widening on LiveCodeBench
To compare the progressive widening with AB-MCTS, we conducted an experiment on LiveCodeBench for progressive widening with various parameters. We used deepseek-v3-0324 with a generation budget $2^{7}$ ($n=5$). The results are shown in Table 4. While an adequate progressive widening parameter leads to strong performance comparable to AB-MCTS, its effectiveness is highly sensitive to hyperparameters. For example, it produces the worst result with $(k,\alpha)=(1,0.45)$. Furthermore, the $(k,\alpha)=(10,0.55)$ setting shows search instability, leading to the highest variance. In contrast, AB-MCTS is robust without such tuning, demonstrating its practical advantage.
Table 5: Comparison of Pass@1 and Pass@2 on ARC-AGI with GPT-4o.
Method Pass@1 Pass@2
Repeated Sampling 14.0 ± 1.7 15.0 ± 1.0
Sequential Refinement 7.7 ± 0.6 8.7 ± 0.9
Standard MCTS 8.0 ± 1.0 9.0 ± 1.5
AB-MCTS-M 11.0 ± 1.0 12.3 ± 1.2
AB-MCTS-A (Gaussian) 13.0 ± 3.6 13.0 ± 3.6
AB-MCTS-A (Beta) 12.7 ± 0.6 14.0 ± 2.1
C.5 AB-MCTS-M vs. AB-MCTS-A: Analysis and Selection
This paper introduces two adaptive branching algorithms: AB-MCTS-M and AB-MCTS-A. This section offers considerations for selecting between them, drawing upon our experimental findings and their distinct underlying mechanisms.
When outcome quality is the main priority, AB-MCTS-M is often the preferred choice because of its consistently strong performance (Table 1 and 2). For node selection, this algorithm uses MCMC, an iterative procedure that improves effectiveness but incurs some computational cost per selection. For applications with strict time constraints, AB-MCTS-A offers a lighter alternative, as its Gaussian and Beta variants employ analytically tractable posterior updates. However, when the LLM-inference time or the time required to evaluate the generated candidate solutions dominate, the extra time spent on MCMC has little impact on total runtime.
While both algorithms feature adaptive search, Figures 5 and 10 show that AB-MCTS-A tends to construct wider search trees. This tendency is attributed to its core design: at each depth, AB-MCTS-A chooses between selecting a GEN node or a CONT node. Reaching depth $d$ requires $d$ consecutive CONT choices, so deeper paths become geometrically less likely than wider expansions (see also Section A.4). Therefore, for tasks such as ARC-AGI where broader exploration is considered particularly beneficial, AB-MCTS-A can be preferable.
Ultimately, the choice depends on the application’s primary goal: AB-MCTS-M often performs well when outcome quality is prioritized, whereas AB-MCTS-A offers advantages in computational efficiency and for tasks that benefit from broader exploration. However, both provide capable adaptive search strategies.
C.6 Pass@1 vs. Pass@2 on ARC-AGI
In Table 1, we reported Pass@2 scores for ARC-AGI to follow the standard evaluation protocol defined by the benchmark [17]. For completeness and transparency, we also report the corresponding Pass@1 results in Table 5. As shown in Table 5, the results indicate consistent relative performance between Pass@1 and Pass@2. In particular, the ranking of methods remains unchanged, confirming that the conclusions presented in the main text are robust to the evaluation metric.
C.7 Analysis of Adaptive Search Behavior via Node Degree Distribution
To gain a granular view of the adaptive nature of our methods, we analyze the node degree distribution of the search trees generated on LiveCodeBench with DeepSeek-V3, illustrated in Figure 12. The results show two distinct behaviors. AB-MCTS-M clearly favors a depth-focused search, with around 90% of its non-leaf nodes having a small degree (1–3). However, its long-tail distribution, with degrees up to 40, confirms that it adaptively broadens the search when required. In contrast, AB-MCTS-A performs a much wider search. Low-degree nodes (1–3) account for only about 30% of its non-leaf nodes, and the distribution spreads broadly, with degrees exceeding 100. This adaptive shaping of the search tree, which differs starkly from the more rigid patterns of baselines, provides strong evidence for the flexibility of our framework.
Figure 12: Node-degree distribution reveals the adaptive search nature of our methods. The plot displays the frequency of nodes (y-axis, log scale) for each degree (x-axis) under a search budget of $N=128$.
C.8 Ablation on the Fixed Branching Factor $w$ in Standard MCTS
Table 6: Performance of Standard MCTS with Different Fixed Branching Factors ($w$) on LiveCodeBench. Pass@1 scores are reported. The best result is shown in bold.
Metric $w=3$ $w=5$ $w=10$
Pass@1 0.429 ± 0.018 0.432 ± 0.021 0.402 ± 0.015
We tested standard MCTS with fixed branching factors ($w$) of 3, 5, and 10 on LiveCodeBench using DeepSeek-V3. The results are summarized in Table 6. We find that $w=5$ achieves the best performance among the tested values. Crucially, performance degrades with both smaller and larger widths, highlighting the sensitivity of the baseline to this hyperparameter. This finding underscores a key advantage of our adaptive method, which dynamically adjusts the effective branching factor during the search.
C.9 Robustness of AB-MCTS Performance
In other baseline methods, the branching factor is either predetermined or determined by hyperparameters. For example, we need to predefine the branching factor in standard MCTS. However, the efficiency of the width vs depth strongly depends on the type of tasks and LLMs. This is reflected in Table 1 and Table 2, where AB-MCTS shows robust performance across various task types, while other methods excel for some tasks, but not for others. This is due to the adaptive branching nature of AB-MCTS, where the algorithm adapts to the wide or deep direction depending on the observed rewards. This is beneficial in the context of LLM inference-time scaling, since recently LLM has been used for solving various kinds of tasks, e.g., math tasks, coding tasks, etc., and a search algorithm that works for various task types out of the box is in high demand.
C.10 Towards Pushing the Pareto Frontier of LLM Inference-Time Scaling
In this section, we propose the following two potential directions for future work aimed at pushing the Pareto frontier of LLM inference-time scaling:
  1. 1.
    Enhanced Adaptivity via Difficulty Estimation: We propose to enhance our algorithm’s existing depth-width balancing by explicitly estimating problem difficulty from collected rewards. For difficult problems (identified by low rewards), the strategy would dynamically switch, e.g., from a deep search (AB-MCTS-M) to a wide one (AB-MCTS-A). This is motivated by findings that optimal search strategies depend on difficulty [27].
  2. 2.
    Collaborative Search with Multiple LLMs: We also propose a method that leverages the diverse strengths of different LLMs within a single search. This is implemented by extending AB-MCTS with multiple GEN nodes, one for each LLM.
We elaborate on the specific methodology and preliminary experimental results for the second direction (Collaborative Search) in Section D.